OpenStudy (grace111):

Will fan and medal whoever helps me . A sequence starts at 500 and 80 is subtracted each time . 500 , 420 , 340 ... The sequence continues in the same way . Write the FIRST TWO NUMBERS in the sequence which are less than zero . Thank you :)

OpenStudy (shiraz14):

This is an A.P. with T1=500 & d=-80. For an A.P., Tn=T1+(n-1)d. So we have: Tn = 500 + (n-1)(-80) < 0 80n > 580 --> n>7.25 n=8 When n=8, T8 = 500 + 7(-80) = 500 - 560 = -60 (Answer) The next term, T9 = -60 - 80 = -140 (Answer)